A rocket, initially at rest on the ground, accelerates straight upward from rest with constant acceleration 44.1 m/s2 . The acceleration period lasts for time 8.00 s until the fuel is exhausted. After that, the rocket is in free fall.
Part A
Find the maximum height ymax reached by the rocket. Ignore air resistance and assume a constant acceleration due to gravity equal to 9.80 m/s2
1
Expert's answer
2019-10-07T11:03:52-0400

Initial speed:


"v=at"

Initial height:


"h=0.5at^2"

From the conservation of energy:


"mgH=mgh+0.5mv^2\to H=gh+0.5v^2"

"H=0.5agt^2+0.5(at)^2=0.5at^2(a+g)"

"H=0.5(44.1)(8)^2(44.1+9.8)=76100\ m=76.1\ km"